Write an equation for an ellipse with major axis of length 12 and foci at(7,2) and (1,2).
Answer by lwsshak3(11628)

Write an equation for an ellipse with major axis of length 12 and foci at(7,2) and (1,2).
**
Based on given foci, equation is that of an ellipse with horizontal major axis: Its standard form:
(x-h)^2/a^2+(y-k)^2/b^2=1
For given equation:
center: (4,2)
given length of major axis=12=2a
a=6
a^2=36
..
c=4 (distance from center to focus on horizontal major axis)
c^2=a^2-b^2
b^2=a^2-c^2=36-16=20
Equation of given ellipse:
(x-4)^2/36+(y-2)^2/20=1